OpenStudy (anonymous):

What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x

OpenStudy (anonymous):

can someone plz help me

OpenStudy (jdoe0001):

do you have your trig formulas sheet?

OpenStudy (anonymous):

no actually this is suppose to be Algebra II.... This is hard

OpenStudy (anonymous):

are uu talking about the sin, cos, tan sheet

OpenStudy (jdoe0001):

mmmm, that looks like trigonometry to me

OpenStudy (jdoe0001):

sin(x+p) is what I'm reading, did you mean to say \[sin(x+\pi)\]?

OpenStudy (anonymous):

on my exam is reads the way I posted it. With the answer choices. My teacher tell s to use rules for a simple periodic function. She has me lost

OpenStudy (anonymous):

if it was suppose to be the way u wrote it is the answer in my selection

OpenStudy (jdoe0001):

do you see in your "sine, cos, tan sheet" something like -> http://upload.wikimedia.org/math/a/9/6/a96a80bf599328734b91591338136fcc.png

OpenStudy (anonymous):

yes that was in the book

OpenStudy (jdoe0001):

ok, then use the \[sin(\alpha+\beta)\] and work with "pi" no "p" which is the only way you'd get an answer :)

OpenStudy (anonymous):

so that is -sin(x)

OpenStudy (jdoe0001):

well, how did you get that?

OpenStudy (anonymous):

guessing... but me trying to work it out now I say it would look like sin(x)+sin(pi) but Im stuck

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

lol

OpenStudy (jdoe0001):

ok, let me start a bit :)

OpenStudy (jdoe0001):

\[sin(x+\pi) = sin(x)*cos(\pi)+cos(x)sin(\pi)\] now, WHAT is the cos(\pi) and sin(\pi)? check your "Unit Circle" if you have to :)

OpenStudy (jdoe0001):

cos(pi) and sin(pi) I meant :)

OpenStudy (anonymous):

ok its -1, 0

OpenStudy (jdoe0001):

ok, forgetting the "x" sin/cos values, use THOSE numbers in the "pi" sections :)

OpenStudy (jdoe0001):

keep in mind that, when you multiply by "0", no matter whatever factor, it'd yeild "0" :)

OpenStudy (anonymous):

ooo so its sin(x)

OpenStudy (anonymous):

I say that because -1 times 0 is zero

OpenStudy (anonymous):

hold on let me try again had to reread what u said do

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

ok I really don't wanna sound stupid but I can't figure this out. Can u show me a set up or give a a hint. I really need to learn this. :(

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

\[sin(x+\pi) = sin(x)*cos(\pi)+cos(x)sin(\pi)= [sin(x)*-1]+[cos(x)*0]\] now multiply those :)

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

I get zero

OpenStudy (jdoe0001):

from -> (sin(x)*-1(+(cos(x)*0)?

OpenStudy (jdoe0001):

let's see sin(x)*-1 + (cos(x)*0) = ?

OpenStudy (jdoe0001):

remember, order of operations :)

OpenStudy (anonymous):

would it be -sin(x) because -1 times x is -1 and 0 times x is 0

OpenStudy (jdoe0001):

yes, because -1 times ANYTHING is -1(ANYTHING), and 0 times ANYTHING is 0

OpenStudy (anonymous):

but when I choice that on as my answer I got it wrong. I'm with James Madison High school form Ashworth College. they dont help at all thanks. are u will to help another day?

OpenStudy (anonymous):

sorry about he misspelled words

OpenStudy (jdoe0001):

sure